OpenStudy (anonymous):

A 6.5 L sample of oxygen at 60C and 0.76 atm is allowed to expand to 13.0 L. The tem- perature remains constant. What is the final pressure?

OpenStudy (anonymous):

Dear Sandy, Boyles Law applies for gases at a constant temperature, and amount (mole). Boyles Law: \[P_1V_1=P_2V_2\]\[\frac{ P_1V_1 }{ V_2 }=P_2\] Problem: Initial volume (V1): 6.5 L Final volume (V2): 13.0 L Initial Pressure (P1): 0.76 atm Final Pressure (P2): ? Temperature (C): 60 C + 273.15 = 333 K \[\frac{ (0.76 atm * 6.5 L) }{ 13.0 L } = 0.38 atm\] The final pressure is 0.38 atm. If you have any questions please do not hesitate to email! Sincerely, N